7. An ice skater can be modeled bv a cylinder of radius R and length L (trunk and legs together), 2 small thin cylinders. of radius r and length l (arms), and a sphere also of radius R (head). The head has mass mh, the trunk mt, and the arms ma. The ice skater rotates about her central axis. Find the moment of inertia of the ice skater with arms outstretched horizontally and then with arms vertical.
Answer #1

We know that the moment of inertia of cylinder is 1/2 MR2 and moment of inertia of sphere is 2/5 MR2 and the moment of inertia of spherical surface  is  2/3 MR2 where M is the mass and R is the radius.

Using this formula, we can solve this question.

For the cylinder whose radius is R and length is L and mass is M, the moment of inertia will be = 1/2 MR2. Similarly , head has mass mh and trunk has mass mt , the moment of inertia will be 1/2 mhR2 + 1/2 mtR2 .

If you add them , we will get the below results.

Ihorizontal = 2/5mhR2 + 1/2 mtR2 + maR2 + maR2 = (2/5 mh + 1/2 mt + 2ma) R2 + 2/3 mal2 + 1/2 mar2 + 2maRL

sililary you can get the value of Ivertical. . Here the area of arms will be 3maR2 + 4maRR

(2/5 mh + 1/2 mt + 2ma) R2 + 3maR2 + 4maRR.

Hence , the result.